OpenStudy (anonymous):

−2x − 13 = 8x + 7, which order of operations completely solves for x

OpenStudy (anonymous):

Division, Addition, and Subtraction. -2x - 13 = 8x + 7 We need to isolate the variable on one side. So I'm going to ADD 2x to both sides. -2x + 2x - 13 = 8x + 7 + 2x -13 = 10x + 7 SUBTRACT 7 from both sides. -13 - 7 = 10x + 7 -7 -20 = 10x DIVIDE 10 by both sides. -20/10 = 10x/10 x = -2 Check your answer. -2x - 13 = 8x + 7 -2(-2) - 13 = 8(-2) + 7 4 - 13 = -16 + 7 -9 = -9 Correct.

OpenStudy (anonymous):

thank you